Family incomes have a mean of $60 000 with a standard deviation of $20 000. The data is normally distributed. What is the probability of a randomly chosen family having an income greater than $50 000? Provide evidence of your work by typing out your full solution.

ASP


1
Expert's answer
2021-08-19T13:19:18-0400

"\mu=60000 \\
\sigma=20000 \\
P(X>50000) = 1-P(X<50000) \\
= 1- P(Z< \frac{50000-60000}{20000}) \\
= 1 -P(Z< -0.5) \\
=1 -0.3085 \\
= 0.6915"